# Variationality of Conformal Geodesics in dimension 3

### v1

@@ 1. Conformal geodesics
Following Bailey-Eastwood [3] we split (1) into its tangential part
@@
${}^{\circ}.$ Equation (1) is covariant, i.e. it does not depend on a choice of coordinates on $M$, and so transforms naturally under diffeomorphisms. Next, we observe that a change of a representative $g$ of the conformal structure $[g]$ leads to the same differential equation. Consequently, (1) is conformally invariant: if $\phi:(M,g)\to(\bar{M},\bar{g})$ is a conformal diffeomorphism, then it transforms the equation for conformal geodesics of $g$ to that of $\bar{g}$. Equivalently, $\phi$ maps conformal geodesics of $g$ to those of $\bar{g}$.
@@
A constrained version of the parametrized conformal geodesic equation is based on normalization of the length parameter $|u|^{2}=1$ (depending on a metric $g$ in the conformal class; in pseudo-Riemannian case this should be changed to ±$\pm 1$), which implies $u\cdot a=0$ and furthermore $u\cdot\nabla_{u}a+|a|^{2}=0$. Combining this with (2) yields $P(u,u)=-\frac{1}{2}|a|^{2}$ and hence (1) with such constraints can be written as
@@
[…]
@@ 2. The inverse variational problem
In the case of second-order $k=2$ the following is known. In dimension $n+1=2$ every path structure is variational as was demonstrated by N. Sonin in 1886 [20]. Dimension $n+1=3$ was investigated in detail by Davis and Duglas [7, 9]. Higher dimensions were also investigated in the literature, and indeed a generic ODE is not variational; we refer to [16] for a discussion.
@@
Let a reparametrization invariant ODE system of 3rd order be variational. Then it is variational in the class of Lagrangians of order 2: there exists a function $L=L({\boldsymbol{x}},\dot{{\boldsymbol{x}}},\ddot{{\boldsymbol{x}}})$ homogeneous of total degree 11 11 and affine in 2-jets $\ddot{{\boldsymbol{x}}}$, whose extremals are precisely the curves of the ODE.
@@
By the Vainberg–Tonti formula [17], we may assume without loss of generality that the Lagrangian of the variational 3rd order system is of 3rd order: $\hat{L}=\hat{L}({\boldsymbol{x}},\dot{{\boldsymbol{x}}},\ddot{{\boldsymbol{x}}},\dddot{{\boldsymbol{x}}})$. In the corresponding Euler–Lagrange equation $\mathop{\rm EL}\nolimits(\hat{L})$ the 6th order terms come from $\frac{d^{3}}{dt^{3}}\frac{\partial\hat{L}}{\partial\dddot{x}^{i}}$. Hence the Lagrangian has the following form:
@@
Vanishing of the 5th order jets in $\mathop{\rm EL}\nolimits(\hat{L})$ is equivalent to the system
@@
which is of the 2nd order. Furthermore, since the 4th order terms in $\mathop{\rm EL}\nolimits(\tilde{L})$ vanish, the Lagrangian $\tilde{L}$ must be affine in $\ddot{{\boldsymbol{x}}}$. Finally, since by our assumptions a reparameterization of a solution ${\boldsymbol{x}}(t)\mapsto{\boldsymbol{x}}(\tau(t))$ is also a solution, the Lagrangian $\tilde{L}$ is homogeneous of total degree 1. ∎
@@ 3. The flat and homogeneous cases
The meaning of $L$ in arc-length parameter is the classical torsion $\tau$ of space curves (more general Lagrangians $L=L(\kappa,\tau)$ depending on curvature and torsion were considered in [13]). The Lagrangian $L$ is invariant with respect to conformal rescaling (the same for all metrics $g=\lambda\cdot ds^{2}_{\text{Eucl}}$) but is not covariant and hence not conformally invariant. Also note that the action
[…]
@@ 4. The general Lagrangian
$\kappa\tau=g(b,B)=g(P^{\sharp}u,B)=P(T,B)=-\mathop{\rm Ric}\nolimits(T,B).$
@@
Thus, we conclude that $L$ is covariant and isometry (but not conformally) invariant, and that the corresponding action $\mathcal{L}$ is projectively invariant. Thus extremals are reparametrization invariant curves.
@@ 5. The proof
Theorem 1 is a curved version of Theorem 2; we will prove it in steps following the strategy from Section 3.
@@
which obviously implies $\mathop{\rm ord(EL)}\nolimits<6$.
@@
$\displaystyle$==  $\displaystyle\,2\mathop{\rm alt}\nolimits_{[ij]}\Bigl{(}\frac{\ell}{A^{2}}\sqrt{|g|}\epsilon_{kij}u^{k}-\frac{2\ell}{A^{4}}\Bigl{(}\ell^{2}a_{i}-(u\cdot a)u_{i}\Bigr{)}\sqrt{|g|}\epsilon_{pqj}u^{p}a^{q}\Bigr{)}.\quad\qquad$ (14)
@@ 5. At last, again using (14), let us compute
$\begin{split}\Bigl{(}2\frac{\partial^{3}L}{\partial x^{d}\partial\ddot{x}^{i}\partial\dddot{x}^{j}}&-3\frac{\partial^{3}L}{\partial x^{d}\partial\ddot{x}^{j}\partial\dddot{x}^{i}}\Bigr{)}u^{d}-\Bigl{(}2\frac{\partial^{3}L}{\partial\ddot{x}^{r}\partial\ddot{x}^{i}\partial\dddot{x}^{j}}-3\frac{\partial^{3}L}{\partial\ddot{x}^{r}\partial\ddot{x}^{j}\partial\dddot{x}^{i}})\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}u^{d}\\ &\doteq\left(-\frac{10}{|a|^{4}}\epsilon_{lij}u^{l}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}+\frac{16}{|a|^{6}}\epsilon_{pqj}u^{p}a_{i}a^{q}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}\right.\\ &-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}\partial_{d}(\Gamma_{bc}^{q})u^{b}u^{c}a_{i}+\frac{4}{|a|^{4}}\epsilon_{pqj}u_{i}u^{p}a^{q}u_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}-\frac{24}{|a|^{6}}\epsilon_{pqi}u^{p}a_{j}a^{q}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}\\ &\left.+\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}+\frac{6}{|a|^{4}}\epsilon_{pri}u^{p}\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}a_{j}-\frac{6}{|a|^{4}}\epsilon_{pqi}u_{j}u^{p}a^{q}u_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}\right)u^{d}\\ &-\left(-\frac{10}{|a|^{4}}\epsilon_{lij}u^{l}a_{r}-\frac{4}{|a|^{4}}\epsilon_{prj}u^{p}a^{i}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}\delta_{ir}a^{q}+\frac{4}{|a|^{4}}\epsilon_{pqj}u_{i}u_{r}u^{p}a^{q}+\frac{16}{|a|^{6}}\epsilon_{pqj}a_{i}u^{p}a^{q}a_{r}\right.\\ &\left.+\frac{6}{|a|^{4}}\epsilon_{pri}u^{p}a_{j}+\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}\delta_{jr}a^{q}-\frac{6}{|a|^{4}}\epsilon_{pqi}u_{j}u_{r}u^{p}a^{q}-\frac{24}{|a|^{6}}\epsilon_{pqi}a_{j}u^{p}a^{q}a_{r}\right)\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}u^{d}\\ &=\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}u^{d}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}u^{d}-\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}u^{d}\\ &+\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}u^{d}=0.\end{split}$ (40)
@@ 6. On conformal invariance
In [10] the Lagrangian $L=\frac{u\cdot E}{|u|^{2}}$ of order 3 was considered, where $E$ is equal to $b=\nabla_{u}a$ minus the expression in the rhs of (1) (in other words, components of the vector $E$ give the equations of parametrized conformal geodesics). This Lagrangian is conformally invariant, but again its extremals are more plentiful than just conformal geodesics; see also [21] for similar results.
@@
[…]
@@
¯$\bar{\nabla}_{X}Y=\nabla_{X}Y-X(\varphi)Y-Y(\varphi)X+g(X,Y)\nabla^{g}\varphi.$
@@
$\displaystyle\bar{b}$ =¯$\displaystyle=\bar{\nabla}_{\bar{u}}\bar{a}=e^{\varphi}\bigl{(}\nabla_{u}\bar{a}-\partial_{u}(\varphi)\bar{a}-\partial_{\bar{a}}(\varphi)u\bigr{)}$
@@
In order to verifify (43) we choose geodesic coordinates for $g$ at the point $x_{0}\in\gamma\subset M$, where all expressions simplify and the identiy follows by a simple straightforward computation. ∎
Let us conclude by mentioning that the problem of existence of (strictly) conformally invariant Lagrangian for (1) is is still open. If such exists, it must be sufficiently degenerate and possibly of higher order. It may be interesting to see if the invariant variational bicomplex technique [15, 2] would work in this problem.

### v2

@@ 1. Conformal geodesics
Following Bailey–Eastwood [3] we split (1) into its tangential part
@@
[…]
@@
A constrained version of the parametrized conformal geodesic equation is based on normalization of the length parameter $|u|^{2}=1$ (depending on a metric $g$ in the conformal class; in pseudo-Riemannian case this should be changed to $\pm 1$), which implies $u\cdot a=0$ and furthermore $u\cdot\nabla_{u}a+|a|^{2}=0$. Combining this with (2) yields $P(u,u)=-\frac{1}{2}|a|^{2}$ and hence (1) with such constraints can be written as
@@
The equation for unparametrized conformal geodesics is variational with Lagrangian $L=V/A^{2}$, where $V=\mathop{\rm Vol}\nolimits_{g}(u,a,\nabla_{u}a)$ and $A=\mathop{\rm Area}\nolimits_{g}(u,a)$; see (12). The corresponding action is $\int L\,dt=\int\tau\,ds$, where $\tau$ is the torsion of the curve and $s$ is the arc length parameter.
@@ 2. The inverse variational problem
In the case of second-order $k=2$ the following is known. In dimension $n+1=2$ every path structure is variational as was demonstrated by N. Sonin in 1886 [27]. Dimension $n+1=3$ was investigated in detail by Davis and Douglas [9, 11]. Higher dimensions were also investigated in the literature, and indeed a generic ODE is not variational; we refer to [20] for a discussion.
@@
Let a reparametrization invariant ODE system of 3rd order be variational. Then it is variational in the class of Lagrangians of order 2: there exists a function $L=L({\boldsymbol{x}},\dot{{\boldsymbol{x}}},\ddot{{\boldsymbol{x}}})$ homogeneous of total degree $1$ and affine in 2-jets $\ddot{{\boldsymbol{x}}}$, whose extremals are precisely the curves of the ODE.
@@
[…]
@@
[…]
@@
[…]
@@ 3. The flat and homogeneous cases
The meaning of $L$ in the arc length parameter is the classical torsion $\tau$ of space curves (more general Lagrangians $L=L(\kappa,\tau)$ depending on curvature and torsion were considered in [15]). The Lagrangian $L$ is invariant with respect to conformal rescaling (the same for all metrics $g=\lambda\cdot ds^{2}_{\text{Eucl}}$) but is not covariant (it depends on Euclidean coordinates) and hence is not conformally invariant. Also note that the action
$\mathcal{L}(\gamma)=\int_{\gamma}L\,dt$ (10)
@@ 4. The general Lagrangian
$\kappa\tau=g(b,B)=g(P^{\sharp}u,B)=P(T,B)=\mathop{\rm Ric}\nolimits(T,B).$
@@
Thus, we conclude that $L$ is covariant (coordinate-independent due to use of covariant derivative) and isometry (but not conformally) invariant, and that the corresponding action $\mathcal{L}$ is projectively invariant. Thus extremals are reparametrization invariant curves.
@@ 5. The proof
Theorem 1 is a curved version of Theorem 2; we will prove it in steps following the strategy from Section 3. The computation is demanding; we use several tricks to simplify it. We first check that the Euler–Lagrange equation has order at most 5, then subsequetly reduce this to 4 and 3. Finally we show it has the same rank in $\dddot{x}$ as the equation of conformal geodesics and vanishes modulo this equation.
@@
which obviously implies $\mathop{\rm ord(EL)}\nolimits<6$ because $L$ is affine in $\dddot{x}$.
@@
$\displaystyle=$  $\displaystyle\,2\mathop{\rm alt}\nolimits_{[ij]}\Bigl{(}\frac{\ell}{A^{2}}\sqrt{|g|}\epsilon_{kij}u^{k}-\frac{2\ell}{A^{4}}\Bigl{(}\ell^{2}a_{i}-(u\cdot a)u_{i}\Bigr{)}\sqrt{|g|}\epsilon_{pqj}u^{p}a^{q}\Bigr{)}.\quad\qquad$ (14)
@@ 5. At last, again using (14), let us compute
$\begin{split}\Bigl{(}2\frac{\partial^{3}L}{\partial x^{d}\partial\ddot{x}^{i}\partial\dddot{x}^{j}}&-3\frac{\partial^{3}L}{\partial x^{d}\partial\ddot{x}^{j}\partial\dddot{x}^{i}}\Bigr{)}u^{d}-\Bigl{(}2\frac{\partial^{3}L}{\partial\ddot{x}^{r}\partial\ddot{x}^{i}\partial\dddot{x}^{j}}-3\frac{\partial^{3}L}{\partial\ddot{x}^{r}\partial\ddot{x}^{j}\partial\dddot{x}^{i}})\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}u^{d}\\ &\hskip-15.0pt\doteq\left(-\frac{10}{|a|^{4}}\epsilon_{lij}u^{l}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}+\frac{16}{|a|^{6}}\epsilon_{pqj}u^{p}a_{i}a^{q}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}\right.\\ &\hskip-15.0pt-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}\partial_{d}(\Gamma_{bc}^{q})u^{b}u^{c}a_{i}+\frac{4}{|a|^{4}}\epsilon_{pqj}u_{i}u^{p}a^{q}u_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}-\frac{24}{|a|^{6}}\epsilon_{pqi}u^{p}a_{j}a^{q}a_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}\\ &\hskip-15.0pt\left.+\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}+\frac{6}{|a|^{4}}\epsilon_{pri}u^{p}\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}a_{j}-\frac{6}{|a|^{4}}\epsilon_{pqi}u_{j}u^{p}a^{q}u_{r}\partial_{d}(\Gamma^{r}_{bc})u^{b}u^{c}\right)u^{d}\\ &\hskip-15.0pt-\left(-\frac{10}{|a|^{4}}\epsilon_{lij}u^{l}a_{r}-\frac{4}{|a|^{4}}\epsilon_{prj}u^{p}a^{i}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}\delta_{ir}a^{q}+\frac{4}{|a|^{4}}\epsilon_{pqj}u_{i}u_{r}u^{p}a^{q}+\frac{16}{|a|^{6}}\epsilon_{pqj}a_{i}u^{p}a^{q}a_{r}\right.\\ &\hskip-15.0pt\left.+\frac{6}{|a|^{4}}\epsilon_{pri}u^{p}a_{j}+\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}\delta_{jr}a^{q}-\frac{6}{|a|^{4}}\epsilon_{pqi}u_{j}u_{r}u^{p}a^{q}-\frac{24}{|a|^{6}}\epsilon_{pqi}a_{j}u^{p}a^{q}a_{r}\right)\partial_{d}(\Gamma_{bc}^{r})u^{b}u^{c}u^{d}\\ &\hskip-15.0pt=\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}u^{d}-\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}u^{d}-\frac{6}{|a|^{4}}\epsilon_{pqi}u^{p}a^{q}\partial_{d}(\Gamma_{bcj})u^{b}u^{c}u^{d}\\ &\hskip-15.0pt+\frac{4}{|a|^{4}}\epsilon_{pqj}u^{p}a^{q}\partial_{d}(\Gamma_{bci})u^{b}u^{c}u^{d}=0.\end{split}$ (40)
@@ 6. On conformal invariance
In [12] the Lagrangian $L=\frac{u\cdot E}{|u|^{2}}$ of order 3 was considered, where $E$ is equal to $b=\nabla_{u}a$ minus the expression in the right-hand side of (1) (in other words, components of the vector $E$ give the equations of parametrized conformal geodesics). This Lagrangian is conformally invariant, but again its extremals are more plentiful than just conformal geodesics; see also [28] for similar results.
@@
This conformal invariance is equivalent to the claim that $\int L\,dt$ has identically same Euler–Lagrange equations as its conformally related action $\int\bar{L}\,dt$. Invariance of finite torsion functional was proved by other methods in [24, Proposition 5.11]. Our result contains a precise formula for the transformation of the Lagrangian, and as a by-product the transformation of the torsion under conformal change of the metric.
@@
$\bar{\nabla}_{X}Y=\nabla_{X}Y-X(\varphi)Y-Y(\varphi)X+g(X,Y)\nabla^{g}\varphi.$
@@
$\displaystyle\bar{b}$ $\displaystyle=\bar{\nabla}_{\bar{u}}\bar{a}=e^{\varphi}\bigl{(}\nabla_{u}\bar{a}-\partial_{u}(\varphi)\bar{a}-\partial_{\bar{a}}(\varphi)u\bigr{)}$
@@
In order to verify (43) we choose geodesic coordinates for $g$ at the point $x_{0}\in\gamma\subset M$, where all expressions simplify and the identiy follows by a straightforward computation. ∎
Denoting by $\measuredangle(v,w)=\mathop{\rm arccos}\nolimits\Bigl{(}\frac{g(v,w)}{|v|_{g}\cdot|w|_{g}}\Bigr{)}$ the angle between vectors $v$ and $w$, we derive from (43) the following relation for any vector $v$ (in fact only the direction of $v$ is important in this formula):
$\bar{L}-\frac{d}{dt}\measuredangle(\pi(\bar{a}),v)=L-\frac{d}{dt}\measuredangle(\pi(a),v).$ (44)
The theory of global differential invariants [19] implies the existence of conformally invariant vector field $\xi$ depending on 4-jet of the curve, complementary to the direction of $u$. A particular form of $\xi$ is not important here and so is omitted, for instance we refer to the conformal Frenet frames in [16].
With such a choice, due to (44) the following Lagrangian has the same extremals as $L$ and (the corresponding form $L\,dt$) is fully conformally invariant:
$\hat{L}=L-\frac{d}{dt}\measuredangle(\pi(a),\xi).$ (45)
Reference [25] operates with similar ideas, but [23] questions its “magical identity involving a globally defined angle”. We interpret this angle as our $\measuredangle(\cdot,\xi)$.
The conformally invariant Lagrangian form given by (45) was already discovered in the conformally flat case as the conformal torsion $T$ times the conformal arc length element $\omega$, see [25, 7] (in other sources, $T$ is denoted by $k_{2}$ and $\mu_{2}$). The corresponding functional (taken with values in $S^{1}={\mathbb{R}}\,\mathop{\rm mod}\nolimits 2\pi$)
$\int T\,\omega=\int\hat{L}\,dt=\int L\,dt=\int\tau\,ds$
is known as the total twist or the Banchoff–White functional (its conformal invariance was proven in [4]). The expressions for the conformal torsion $T$ and the arc length element $\omega$ through the metric curvature $\kappa$, the torsion $\tau$ and the natural parameter $ds$ of a curve in the Euclidean space ${\mathbb{R}}^{3}$ are the following:
$T=\Bigl{(}2\kappa_{s}^{2}\tau+\kappa\,\kappa_{s}\tau_{s}-\kappa\,\kappa_{ss}\tau+\kappa^{2}\tau^{3}\Bigr{)}\Bigl{/}\sqrt[4]{(\kappa_{s}^{2}+\kappa^{2}\tau^{2})^{5}},\quad\omega=\sqrt[4]{\kappa_{s}^{2}+\kappa^{2}\tau^{2}}\,ds.$
Thus we conclude that raising the order of Lagrangian to 4 yields a possibility to choose a conformally invariant Lagrangian $\hat{L}$ for conformal geodesics. Apparently such a Lagrangian $L$ cannot exist in order 3.
Remark 4.
The nomenclature “conformal geodesics” was used differently in [25, 23], namely as the Euler–Lagrange equation for the conformal arc length $\int\omega$. This is a sixth order equation, and it is variational by definition. We however keep using the terminolgy adapted in [30, 17, 3].
Let us conclude by mentioning that it may be interesting to see if the invariant variational bicomplex technique [18, 2] would work for the invariant variational problem $\int\hat{L}dt=\int T\,\omega$. This question will be addressed in the forthcoming paper [21].